# Smooth Tests for Normality in ANOVA

### v3

@@ 3.1 Same mean and same variance
$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}\left(\hat{Z}_{ij}\right)=$ $\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left\{\left(\pi_{k}(Z_{ij})-\mathrm{E}\left[\pi_{k}(Z)\right]\right)-d_{1k}e_{ij}-\frac{d_{2k}}{2}\left(e_{ij}^{2}-1\right)\right\}$ (3.5)
$+\mathrm{E}\left[\pi_{k}(Z)\right]+o_{p}\left(\frac{1}{\sqrt{N}}\right),$
@@
$\sqrt{N}\left(\hat{\Psi}_{K}^{2}-\bm{a}_{K}^{\top}\bm{\Sigma}_{K}^{-1}\bm{a}_{K}\right)\to_{d}\mathcal{N}\left(0,4\bm{a}_{K}^{\top}\bm{\Sigma}_{K}^{-1}\bm{\Xi}_{K}\bm{\Sigma}_{K}^{-1}\bm{a}_{K}\right),$ (3.6)
where $\bm{a}_{K}\equiv(a_{1},\ldots,a_{K})^{\top}=(\mathrm{E}[\pi_{1}(Z)],\ldots,\mathrm{E}[\pi_{K}(Z)])^{\top}$ and $\bm{\Xi}_{K}=\left(\xi_{kl}\right)_{K\times K}$ is given by
$\xi_{kl}=$ $\mathrm{E}\left[\pi_{k}(Z)\pi_{l}(Z)\right]-a_{k}a_{l}-\left[d_{1k}d_{3l}+d_{1l}d_{3k}\right]+d_{1k}d_{1l}+\frac{1}{2}\left[a_{k}d_{2l}+a_{l}d_{2k}\right]$
$-\frac{1}{2}\left[d_{2k}d_{4l}+d_{2l}d_{4k}\right]+\frac{1}{2}\left[d_{1k}d_{2l}+d_{1l}d_{2k}\right]\mathrm{E}\left[e^{3}\right]+\frac{d_{2k}d_{2l}}{4}\left[\mathrm{E}\left[e^{4}\right]-1\right].$
From Theorem 2, it follows that under the alternative hypothesis $H_{1}$ in (3.4), if there exists at least one $1\leq k\leq K$ such that
$a_{k}\equiv\mathrm{E}[\pi_{k}(Z)]=\int_{-\infty}^{\infty}\pi_{k}(\Phi(x))f(x)\mathrm{d}x=\int_{0}^{1}\pi_{k}(z)\frac{f(\Phi^{-1}(z))}{\phi(\Phi^{-1}(z))}\mathrm{d}z\neq 0,$
then the smooth test statistic satisfies $N\hat{\Psi}_{K}^{2}\to\infty$ in probability as $N\to\infty$, implying that the asymptotic power of the test is $1$. Specifically, the power function is given by
@@
The following theorem presents the theoretical properties of $N\hat{\Psi}_{K}^{2}$ under the local alternatives. Specifically, the rate of $\delta_{N}$ tending to 0 as $N\to\infty$ is crucial to the nontrivial local power of the test.
@@
$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}\left(\hat{Z}_{ij}\right)=$ $\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left\{\left(\pi_{k}(Z_{ij})-\mathrm{E}\left[\pi_{k}(Z)\right]\right)-c_{1k}e_{ij}-\frac{c_{2k}}{2}\left(e_{ij}^{2}-1\right)\right\}$ (3.8)
@@
$N\hat{\Psi}_{K}^{2}\to_{d}\chi_{K}^{2}\left(\bm{\Delta}_{K}^{\top}\bm{\Sigma}_{K}^{-1}\bm{\Delta}_{K}\right),$
@@
It is worthwhile to mention that $N^{-1/2}$ is the critical order of magnitude for our test to achieve nontrivial power against the Pitman-type sequence of local alternatives. If $\delta_{N}=o(N^{-1/2})$, it would be impossible to distinguish the alternatives from the null. Under $H_{1L}$ in (3.7), as long as there exists at least one $1\leq k\leq K$ such that $\Delta_{k}\neq 0$, the proposed test statistic $N\hat{\Psi}_{K}^{2}$ attains nontrivial asymptotic power against the local alternatives because the noncentrality parameter $\bm{\Delta}_{K}^{\top}\bm{\Sigma}_{K}^{-1}\bm{\Delta}_{K}>0$.
@@
We emphasize that the aforementioned smooth test is not “consistent” in the strict sense; that is, the power of the test does not necessarily approach $1$ as $N\to\infty$ under all directions of fixed alternatives. Only under alternatives where $\mathrm{E}[\pi_{k}(Z)]\neq 0$ for at least one $1\leq k\leq K$ does the asymptotic power converge to $1$. Otherwise, if $\mathrm{E}[\pi_{k}(Z)]=0$ for all $1\leq k\leq K$ (but $\mathrm{E}[\pi_{K+1}(Z)]\neq 0$ maybe), $\hat{\Psi}_{K}^{2}$ will fail to detect the discrepancy between $F$ and $\Phi$. For example, let $e\sim\mathcal{U}[-\sqrt{3},\sqrt{3}]$ so that $\mathrm{E}[e]=0$ and $\operatorname{Var}[e]=1$, and consider the first-order orthonormal Legendre polynomial $\pi_{1}(z)=\sqrt{3}(2z-1)$. For $Z=\Phi(e)$, we have
$\mathrm{E}[Z]$ $=\frac{1}{2\sqrt{3}}\int_{-\sqrt{3}}^{\sqrt{3}}\Phi(z)\mathrm{d}z$
@@
which implies $\mathrm{E}[\pi_{1}(Z)]=0$. In fact, smooth tests are neither directional nor omnibus; that is, they maintain reasonable power across a broad—but not universal—range of alternatives, and generally exhibit good finite-sample power properties; see Bera et al., (2013) for further discussion.
3.2 Different means and same variance
Next, we consider ANOVA models in which the group means are allowed to differ, that is,
@@ 3.2 Different means and same variance
where $\{\mu_{j}\}_{j=1}^{J}$ denote the potentially distinct means of each group. Let $\hat{\mu}_{j}=N_{j}^{-1}\sum_{i=1}^{N_{j}}Y_{ij}$ and $\hat{\sigma}^{2}=N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(Y_{ij}-\hat{\mu}_{j}\right)^{2}$ be the estimators of $\mu_{j}$ ($j=1,\ldots,J$) and $\sigma^{2}$, respectively. Correspondingly we define $\hat{\varepsilon}_{ij}=Y_{ij}-\hat{\mu}_{j}$, $\hat{e}_{ij}=\hat{\varepsilon}_{ij}/\hat{\sigma}=(Y_{ij}-\hat{\mu}_{j})/\hat{\sigma}$, and $\hat{Z}_{ij}=\Phi(\hat{e}_{ij})$. The following theorem characterizes the properties of $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}(\hat{Z}_{ij})$ under the model (3.9).
@@
$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}\left(\hat{Z}_{ij}\right)=$ $\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left\{\pi_{k}(Z_{ij})-c_{1k}e_{ij}-\frac{c_{2k}}{2}\left(e_{ij}^{2}-1\right)\right\}+o_{p}\left(\frac{1}{\sqrt{N}}\right),$
as $\min\{N_{1},\ldots,N_{J}\}\to\infty$, $J=o(N^{1/2})$ and $\sum_{j=1}^{J}N_{j}^{-1}=o(1)$. Furthermore, for $\hat{\Psi}_{K}^{2}$ defined in (3.3), we have
$N\hat{\Psi}_{K}^{2}\to_{d}\chi_{K}^{2}.$
Theorem 4 parallels Theorem 1 and Corollary 1, except that it requires additional conditions on the group sample sizes and the number of groups. These conditions are imposed for technical reasons and to account for heterogeneity across group means in model (3.9). The intuition behind the conditions is illustrated as follows. First, the sample size of each group, $N_{j}$, must tend to infinity. In model (3.9), where $\mu_{j}$ can be distinct, only the observations within group $j$ can be used to estimate $\mu_{j}$, making $N_{j}\to\infty$ necessary for the consistency of $\hat{\mu}_{j}$. Second, the number of groups $J$ is required to be finite or to diverge no faster than $N^{1/2}$, and $\sum_{j=1}^{J}N_{j}^{-1}=o(1)$ must hold jointly for $\{N_{j}\}_{j=1}^{J}$ and $J$. This rules out scenarios where the number of groups grows faster than the sample sizes within each group. In contrast, for model (3.1) in the previous section, it suffices that the total sample size $N\to\infty$ to derive asymptotic properties, regardless of the magnitudes of $N_{j}$ and $J$. This is because in model (3.1), all groups share the same mean $\mu$ and variance $\sigma^{2}$, so the observations $\{Y_{ij}\}_{i=1,j=1}^{N_{j},J}$ are i.i.d., allowing them to be pooled for efficient estimation and inference.
The following two theorems establish formally the asymptotic properties of $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}(\hat{Z}_{ij})$ and $\hat{\Psi}_{K}^{2}$ (defined in (3.3)) under the alternatives. The conditions are identical to those in Theorem 4, and the results parallel those in Theorems 2 and 3.
@@
$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}\left(\hat{Z}_{ij}\right)=$ $\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left\{\left(\pi_{k}(Z_{ij})-\mathrm{E}\left[\pi_{k}(Z)\right]\right)-d_{1k}e_{ij}-\frac{d_{2k}}{2}\left(e_{ij}^{2}-1\right)\right\}$
$+\mathrm{E}\left[\pi_{k}(Z)\right]+o_{p}\left(\frac{1}{\sqrt{N}}\right),$
@@
$\sqrt{N}\left(\hat{\Psi}_{K}^{2}-\bm{a}_{K}^{\top}\bm{\Sigma}_{K}^{-1}\bm{a}_{K}\right)\to_{d}\mathcal{N}\left(0,4\bm{a}_{K}^{\top}\bm{\Sigma}_{K}^{-1}\bm{\Xi}_{K}\bm{\Sigma}_{K}^{-1}\bm{a}_{K}\right).$
@@
$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}\left(\hat{Z}_{ij}\right)=$ $\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left\{\left(\pi_{k}(Z_{ij})-\mathrm{E}\left[\pi_{k}(Z)\right]\right)-c_{1k}e_{ij}-\frac{c_{2k}}{2}\left(e_{ij}^{2}-1\right)\right\}$
@@
$N\hat{\Psi}_{K}^{2}\to_{d}\chi_{K}^{2}\left(\bm{\Delta}_{K}^{\top}\bm{\Sigma}_{K}^{-1}\bm{\Delta}_{K}\right).$
@@
3.3 Same mean and different variances
We now extend model (3.1) by allowing the group variances to differ across groups, that is,

### v4

@@ 3.1 Common group means and common group variances
$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}\left(\widehat{Z}_{ij}\right)=$ $\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left\{\left(\pi_{k}(Z_{ij})-\mathbb{E}\left[\pi_{k}(Z)\right]\right)-d_{1k}e_{ij}-\frac{d_{2k}}{2}\left(e_{ij}^{2}-1\right)\right\}$ (3.5)
$+\mathbb{E}\left[\pi_{k}(Z)\right]+o_{p}\left(\frac{1}{\sqrt{N}}\right),$
@@
$\sqrt{N}\left(\widehat{\Psi}_{K}^{2}-\bm{a}_{K}^{\top}\bm{\Sigma}_{K}^{-1}\bm{a}_{K}\right)\overset{d}{\to}\mathcal{N}\left(0,4\bm{a}_{K}^{\top}\bm{\Sigma}_{K}^{-1}\bm{\Xi}_{K}\bm{\Sigma}_{K}^{-1}\bm{a}_{K}\right),$ (3.6)
where $\bm{a}_{K}\equiv(a_{1},\ldots,a_{K})^{\top}=(\mathbb{E}[\pi_{1}(Z)],\ldots,\mathbb{E}[\pi_{K}(Z)])^{\top}$ and $\bm{\Xi}_{K}=\left(\xi_{kl}\right)_{K\times K}$ is given by
$\xi_{kl}=$ $\mathbb{E}\left[\pi_{k}(Z)\pi_{l}(Z)\right]-a_{k}a_{l}-\left[d_{1k}d_{3l}+d_{1l}d_{3k}\right]+d_{1k}d_{1l}+\frac{1}{2}\left[a_{k}d_{2l}+a_{l}d_{2k}\right]$
$-\frac{1}{2}\left[d_{2k}d_{4l}+d_{2l}d_{4k}\right]+\frac{1}{2}\left[d_{1k}d_{2l}+d_{1l}d_{2k}\right]\mathbb{E}\left[e^{3}\right]+\frac{d_{2k}d_{2l}}{4}\left[\mathbb{E}\left[e^{4}\right]-1\right].$
The decomposition (3.5) exhibits different properties from that in Theorem 1. Specifically, the first term on the right-hand side $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}(Z_{ij})$ now has a nonzero expectation under $H_{1}$ given by (3.4), in contrast to its zero mean under $H_{0}$, which constitutes the primary source of power for the proposed test. The latter two terms, $d_{1k}N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}e_{ij}$ and $d_{2k}(2N)^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}(e_{ij}^{2}-1)$, represent the estimation effects of $\widehat{\mu}$ and $\widehat{\sigma}^{2}$ under $H_{1}$, respectively. Moreover, when $H_{0}$ is true, $d_{1k}=c_{1k}$ and $d_{2k}=c_{2k}$, then the decomposition (3.5) coincides with that in Theorem 1.
From Theorem 2, it follows that under $H_{1}$ given by (3.4), if there exists at least one $1\leq k\leq K$ such that
$a_{k}\equiv\mathbb{E}[\pi_{k}(Z)]=\int_{-\infty}^{\infty}\pi_{k}(\Phi(x))f(x)\mathrm{d}x=\int_{0}^{1}\pi_{k}(z)\frac{f(\Phi^{-1}(z))}{\phi(\Phi^{-1}(z))}\mathrm{d}z\neq 0,$
then the smooth test statistic satisfies $N\widehat{\Psi}_{K}^{2}\to\infty$ in probability as $N\to\infty$, implying that the asymptotic power of the test is $1$. Specifically, the power function is given by
@@
The following theorem presents the theoretical properties of $N\widehat{\Psi}_{K}^{2}$ under the local alternatives. Specifically, the rate of $\delta_{N}$ tending to 0 as $N\to\infty$ is crucial to the nontrivial local power of the test.
@@
$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}\left(\widehat{Z}_{ij}\right)=$ $\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left\{\left(\pi_{k}(Z_{ij})-\mathbb{E}\left[\pi_{k}(Z)\right]\right)-c_{1k}e_{ij}-\frac{c_{2k}}{2}\left(e_{ij}^{2}-1\right)\right\}$ (3.8)
@@
$N\widehat{\Psi}_{K}^{2}\overset{d}{\to}\chi_{K}^{2}\left(\bm{\Delta}_{K}^{\top}\bm{\Sigma}_{K}^{-1}\bm{\Delta}_{K}\right),$
@@
The asymptotic decomposition (3.8) is closely related to those obtained under $H_{0}$ and $H_{1}$, while exhibiting a key difference. In particular, compared with the decomposition under $H_{0}$, the leading term $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}(Z_{ij})$ becomes noncentered under $H_{1L}$, with mean shift $\mathbb{E}[\pi_{k}(Z)]=\delta_{N}\Delta_{k}$, which introduces an additional deterministic drift term of order $\delta_{N}$. The two estimation effect terms $c_{1k}N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}e_{ij}$ and $c_{2k}(2N)^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}(e_{ij}^{2}-1)$ coincide with those under $H_{0}$. When $\delta_{N}=N^{-1/2}$, the drift term $\delta_{N}\Delta_{k}$ is of the same order as the stochastic fluctuations, leading to a nondegenerate limit distribution. Consequently, the statistic $N\widehat{\Psi}_{K}^{2}$ converges to a noncentral Chi-square distribution. This also shows that $N^{-1/2}$ is the critical rate for detecting Pitman-type local alternatives: if $\delta_{N}=o(N^{-1/2})$, the alternatives become asymptotically indistinguishable from the null. Under $H_{1L}$ in (3.7), as long as there exists at least one $1\leq k\leq K$ such that $\Delta_{k}\neq 0$, the proposed test statistic $N\widehat{\Psi}_{K}^{2}$ attains nontrivial asymptotic power against the local alternatives because the noncentrality parameter $\bm{\Delta}_{K}^{\top}\bm{\Sigma}_{K}^{-1}\bm{\Delta}_{K}>0$.
@@
We emphasize that the aforementioned smooth test is not “consistent” in the strict sense; that is, the power of the test does not necessarily approach $1$ as $N\to\infty$ under all directions of fixed alternatives. Only under alternatives where $\mathbb{E}[\pi_{k}(Z)]\neq 0$ for at least one $1\leq k\leq K$ does the asymptotic power converge to $1$. Otherwise, if $\mathbb{E}[\pi_{k}(Z)]=0$ for all $1\leq k\leq K$ (but $\mathbb{E}[\pi_{K+1}(Z)]\neq 0$ maybe), $\widehat{\Psi}_{K}^{2}$ will fail to detect the discrepancy between $F$ and $\Phi$. For example, let $e\sim\mathcal{U}[-\sqrt{3},\sqrt{3}]$ so that $\mathbb{E}[e]=0$ and $\operatorname{Var}[e]=1$, and consider the first-order orthonormal Legendre polynomial $\pi_{1}(z)=\sqrt{3}(2z-1)$. For $Z=\Phi(e)$, we have
$\mathbb{E}[Z]$ $=\frac{1}{2\sqrt{3}}\int_{-\sqrt{3}}^{\sqrt{3}}\Phi(z)\mathrm{d}z$
@@
which implies $\mathbb{E}[\pi_{1}(Z)]=0$ and thus no power under this choice. In fact, smooth tests are neither directional nor omnibus; that is, they maintain reasonable power across a broad—but not universal—range of alternatives, and generally exhibit good finite-sample power properties; see bera2013smooth for further discussion.
3.2 Heterogeneous group means with common variances
Next, we consider ANOVA models with heterogeneous group means while maintaining common variances:
@@ 3.2 Heterogeneous group means with common variances
where $\{\mu_{j}\}_{j=1}^{J}$ denote the potentially distinct means of each group. Let $\widehat{\mu}_{j}=N_{j}^{-1}\sum_{i=1}^{N_{j}}Y_{ij}$ and $\widehat{\sigma}^{2}=N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(Y_{ij}-\widehat{\mu}_{j}\right)^{2}$ be the estimators of $\mu_{j}$ ($j=1,\ldots,J$) and $\sigma^{2}$, respectively. Correspondingly we define $\widehat{\varepsilon}_{ij}=Y_{ij}-\widehat{\mu}_{j}$, $\widehat{e}_{ij}=\widehat{\varepsilon}_{ij}/\widehat{\sigma}=(Y_{ij}-\widehat{\mu}_{j})/\widehat{\sigma}$, and $\widehat{Z}_{ij}=\Phi(\widehat{e}_{ij})$. The following theorem characterizes the properties of $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}(\widehat{Z}_{ij})$ under the model (3.9).
@@
$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}\left(\widehat{Z}_{ij}\right)=$ $\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left\{\pi_{k}(Z_{ij})-c_{1k}e_{ij}-\frac{c_{2k}}{2}\left(e_{ij}^{2}-1\right)\right\}+o_{p}\left(\frac{1}{\sqrt{N}}\right),$
as $\min\{N_{1},\ldots,N_{J}\}\to\infty$, $J=o(N^{1/2})$ and $\sum_{j=1}^{J}N_{j}^{-1}=o(1)$. Furthermore, for $\widehat{\Psi}_{K}^{2}$ defined in (3.3), we have
$N\widehat{\Psi}_{K}^{2}\overset{d}{\to}\chi_{K}^{2}.$
Theorem 4 parallels Theorem 1 and Corollary 1, except that it requires additional conditions on the group sample sizes and the number of groups. These conditions are imposed for technical reasons and to account for heterogeneity across group means in model (3.9). The intuition behind the conditions is illustrated as follows. First, the sample size of each group, $N_{j}$, must tend to infinity. In model (3.9), where $\mu_{j}$ can be distinct, only the observations within group $j$ can be used to estimate $\mu_{j}$, making $N_{j}\to\infty$ necessary for the consistency of $\widehat{\mu}_{j}$. Second, the number of groups $J$ is required to be finite or to diverge no faster than $N^{1/2}$, and $\sum_{j=1}^{J}N_{j}^{-1}=o(1)$ must hold jointly for $\{N_{j}\}_{j=1}^{J}$ and $J$. This rules out scenarios where the number of groups grows faster than the sample sizes within each group. In contrast, for model (3.1) in the previous section, it suffices that the total sample size $N\to\infty$ to derive asymptotic properties, regardless of the magnitudes of $N_{j}$ and $J$. This is because in model (3.1), all groups share the common mean $\mu$ and common variance $\sigma^{2}$, so the observations $\{Y_{ij}\}_{i=1,j=1}^{N_{j},J}$ are i.i.d., allowing them to be pooled for efficient estimation and inference.
The following two theorems establish formally the asymptotic properties of $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}(\widehat{Z}_{ij})$ and $\widehat{\Psi}_{K}^{2}$ (defined in (3.3)) under the alternatives. The conditions are identical to those in Theorem 4, and the results parallel those in Theorems 2 and 3.
@@
$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}\left(\widehat{Z}_{ij}\right)=$ $\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left\{\left(\pi_{k}(Z_{ij})-\mathbb{E}\left[\pi_{k}(Z)\right]\right)-d_{1k}e_{ij}-\frac{d_{2k}}{2}\left(e_{ij}^{2}-1\right)\right\}$
$+\mathbb{E}\left[\pi_{k}(Z)\right]+o_{p}\left(\frac{1}{\sqrt{N}}\right),$
@@
$\sqrt{N}\left(\widehat{\Psi}_{K}^{2}-\bm{a}_{K}^{\top}\bm{\Sigma}_{K}^{-1}\bm{a}_{K}\right)\overset{d}{\to}\mathcal{N}\left(0,4\bm{a}_{K}^{\top}\bm{\Sigma}_{K}^{-1}\bm{\Xi}_{K}\bm{\Sigma}_{K}^{-1}\bm{a}_{K}\right).$
@@
$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}\left(\widehat{Z}_{ij}\right)=$ $\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left\{\left(\pi_{k}(Z_{ij})-\mathbb{E}\left[\pi_{k}(Z)\right]\right)-c_{1k}e_{ij}-\frac{c_{2k}}{2}\left(e_{ij}^{2}-1\right)\right\}$
@@
$N\widehat{\Psi}_{K}^{2}\overset{d}{\to}\chi_{K}^{2}\left(\bm{\Delta}_{K}^{\top}\bm{\Sigma}_{K}^{-1}\bm{\Delta}_{K}\right).$
@@
3.3 Common group means with heterogeneous variances
We now extend model (3.1) to allow for heterogeneous group variances, that is,